OpenStudy (anonymous):

Is anyone really good at 8th grade math?

OpenStudy (anonymous):

Yes?

OpenStudy (godlovesme):

question? :)

OpenStudy (anonymous):

(07.01 LC) What is the value of y in the equation 4 + y = –3? (1 point) 7 1 –1 –7 2. (07.01 MC) Mona solved an equation incorrectly, as shown below: Step 1: 5x = 30 Step 2: x = 30 – 5 Step 3: x = 25 Which statement best explains why Step 2 is incorrect in Mona's solution? (1 point) She did not add 5 to 30. She did not divide 30 by 5. She did not multiply 30 by 5. She did not multiply 30 five times. 3. (07.02 LC) How many solutions does the equation –2y + 2y + 3 = 3 have? (1 point) One None Infinitely many Three 4. (07.02 MC) What is the value of x in the equation 4(2x + 14) = 0? (1 point) 9 7 –7 –9 5. (07.02 MC) An equation is shown below: 4(2x – 3) = 1 Which of the following correctly shows the first two steps to solve this equation? (1 point) Step 1: 6x – 7 = 1; Step 2: 6x = 8 Step 1: 6x + 1 = 1; Step 2: 6x = 0 Step 1: 8x – 12 = 1; Step 2: 8x = 13 Step 1: 8x – 3 = 1; Step 2: 8x = 4 6. (07.03 LC) How many solutions does the equation 4y – 4y – 12 = 14 – 2 have? (1 point) One Zero Two Many 7. (07.03 LC) The steps below show the incomplete solution to find the value of x for the equation 8x – 3x – 1 = –2 + 16: Step 1: 8x – 3x – 1 = –2 + 16 Step 2: 8x – 3x – 1 = 14 Step 3: 5x – 1 = 14 Which of these is most likely the next step? (1 point) 5x = 14 5x = 10 5x = 13 5x = 15 8. (07.03 MC) What is the value of y in the equation 2(3y + 7 + 5) = 196 – 16? (1 point) 13 14 26 28 9. (07.03 MC) What is the solution to the following equation? 5(2x – 14) + 23 = 7x – 14 (1 point) 11 14 30 36 10. (07.05 LC) What is the value of a in the equation a = 2 + 3a + 6? (1 point) –8 –4 4 8 11. (07.05 LC) How many solutions does the equation 3x + 6 = –1 – 3 + 4x have? (1 point) Two None One Infinitely many 12. (07.05 MC) Which of these is a simplified form of the equation 9p + 3 = – p + 7 + 2p + 3p? (1 point) 5p = 4 15p = 10 9 = 4 12 = 13 13. (07.05 MC) What is the solution to the equation 4x + 2(x – 3) = 3x + x – 12? (1 point) –3 –1 1 3 14. (07.05 HC) The work of a student to solve the equation 3(3y – 8) = 10 + 3y is shown below: Step 1: 3(3y – 8) = 10 + 3y Step 2: 6y – 5 = 10 + 3y Step 3: 3y – 5 = 10 Step 4: 3y = 15 Step 5: y = 5 In which step did the student first make an error and what is the correct step? (1 point) Step 2: 9y – 24 = 10 + 3y Step 2: 6y + 11 = 10 + 4y Step 3: 6y – 8 = 3(6 + y) Step 3: 3(3y – 8) = 7 + y 15. (07.06 LC) Which statement is true about the equation 4.5z = 4.5z – 4.25 + 6.25? (1 point) It has no solution. It has one solution. It has two solutions. It has infinitely many solutions. 16. (07.06 MC) What is the solution to the equation ? (1 point) n = –2 n = 4 n = n =

OpenStudy (solomonzelman):

I am really poor at math, but I can perhaps try.

OpenStudy (agl202):

wowse!!!!!

OpenStudy (solomonzelman):

1 question at a time, mam. Please !

OpenStudy (anonymous):

flvs?

OpenStudy (solomonzelman):

some online skool :D

Nnesha (nnesha):

flvs = flying vs swimming ???

OpenStudy (agl202):

lol @Nnesha !!!>_>

OpenStudy (anonymous):

so can anyone help me plz?

OpenStudy (solomonzelman):

Nnesha, that is a good topic for a compare-contrast paper:) lol

OpenStudy (solomonzelman):

lets start from question 1, but I am only good at math if you are there to help me (lol). Who know what I will do, but if we do it together, it will be better, wouldn't it?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

What is the value of y in the equation \(\large\color{black}{ 4+y=-3 }\) ? Please, subtract 4 from both sides for me.

OpenStudy (anonymous):

I think #1 is -7

OpenStudy (solomonzelman):

yes, very goot

OpenStudy (solomonzelman):

good*

OpenStudy (anonymous):

because 4+-7=-3

OpenStudy (anonymous):

and I think #2 is b?

OpenStudy (solomonzelman):

Mona solved an equation incorrectly, as shown below: ``` Step 1: 5x = 30 Step 2: x = 30 - 5 Step 3: x = 25 ``` Which statement best explains why Step 2 is incorrect in Mona's solution? (1 point) A. She did not add 5 to 30. B. She did not divide 30 by 5. C. She did not multiply 30 by 5. D. She did not multiply 30 five times. yes, B is correct.

OpenStudy (anonymous):

yay! and for #3 I have no clue

OpenStudy (solomonzelman):

Now, 3. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ How many solutions does the equation \(\large\color{black}{ -2y + 2y + 3 = 3 }\) have? One None Infinitely many Three

OpenStudy (solomonzelman):

it is asking, basically: For how many y-values will the equation above be true ?

OpenStudy (solomonzelman):

you can subtract 2y from sides \(\large\color{black}{ -2y + 3 = -2y+3 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ -2y + 3 = -2y+3 }\) is this statement true when y=1? when y=2? when y=(any number) ?

OpenStudy (anonymous):

that would leave 3=3

OpenStudy (anonymous):

if you subtract -2y from both sides?

OpenStudy (solomonzelman):

the statement (initial statement) would leave 3=3

OpenStudy (solomonzelman):

if you just simplify the initial statement

OpenStudy (solomonzelman):

then regardless of what value is, we will still be having 3=3

OpenStudy (solomonzelman):

so the answer for #3 is ?

OpenStudy (solomonzelman):

what do you mean by 3!?

OpenStudy (solomonzelman):

3! means 1 times 2 times 3 in math (it is a factorial)

OpenStudy (solomonzelman):

if you said that it was option 3, then it is corect

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

What is the value of x in the equation 4(2x + 14) = 0? (1 point) 9 7 -7 -9

OpenStudy (anonymous):

so C??????????????????? for #3

OpenStudy (solomonzelman):

yes, for #3

OpenStudy (anonymous):

okay thats what i thought yay

OpenStudy (solomonzelman):

What is the value of x in the equation 4(2x + 14) = 0? (1 point) 4(2x + 14) = 0 steps 1. divide both sides by 4 2. subtract 14 from both sides 3. divide both sides by 2

OpenStudy (anonymous):

so dividing 4 by 2 and 14?

OpenStudy (anonymous):

?

OpenStudy (solomonzelman):

0 / 4 = ?

OpenStudy (solomonzelman):

4(2x + 14) \ 4 (cancels the 4's), and you get = ?

OpenStudy (anonymous):

0/4=0

OpenStudy (anonymous):

@SolomonZelman